# Forward and Backward State Abstractions for Off-policy Evaluation

### v1

@@ 3 Proposed state abstractions for policy evaluation
Here, we propose model-free (Section 3.1) and model-based irrelevance conditions (Section 3.2) for OPE, and analyze the OPE estimators under these conditions (Theorem 1, Theorem 2, Theorem 3). Motivated by this analysis, we propose our two-step procedure (Section 3.3).
@@ 3.1 Model-free irrelevance conditions
Based on the aforementioned definitions, we can immediately state the following theorem:
Theorem 1 (OPE under model-free irrelevance conditions)
Under $Q^{\pi}$-, $\rho^{\pi}$- or $w^{\pi}$-irrelevance, the corresponding methods remain valid when applied to the abstract state space:
  • Under $Q^{\pi}$-irrelevance, the Q-function-based method (Method 1) remains valid, i.e., the Q-function $Q_{\phi}^{\pi}$ defined on the abstract state space satisfies $\mathbb{E}[f_{1}(Q^{\pi})]=\mathbb{E}[f_{1}(Q^{\pi}_{\phi})]$;
  • Under $\rho^{\pi}$-irrelevance, SIS (Method 2) remains valid, i.e., the IS ratio $\rho_{\phi}^{\pi}$ defined on the abstract state space satisfies $\mathbb{E}[f_{2}(\rho^{\pi})]=\mathbb{E}[f_{2}(\rho^{\pi}_{\phi})]$;
  • Under $w^{\pi}$-irrelevance, MIS (Method 3) remains valid, i.e., the MIS ratio $w_{\phi}^{\pi}$ defined on the abstract state space satisfies $\mathbb{E}[f_{3}(w^{\pi})]=\mathbb{E}[f_{3}(w^{\pi}_{\phi})]$.
Moreover, when $\phi$ satisfies either $Q^{\pi}$-irrelevance or $w^{\pi}$-irrelevance, DRL (Method 4) remains valid, i.e., $Q_{\phi}^{\pi}$ and $w_{\phi}^{\pi}$ defined on the abstract state space satisfy $\mathbb{E}[f_{4}(Q^{\pi},w^{\pi})]=\mathbb{E}[f_{4}(Q^{\pi}_{\phi},w^{\pi}_{\phi})]$.
Theorem 1 validates the four OPE methods presented in Section 2.3 when applied to the abstract state space, under the corresponding irrelevance conditions. Notably, DRL requires weaker irrelevance conditions compared to the Q-function-based method and MIS, owing to its inherent double robustness property. Nevertheless, methods for deriving abstractions that satisfy these conditions (particularly $Q^{\pi}$- and $w^{\pi}$-irrelevance) remain unclear. Furthermore, the state-action-reward triplets transformed via these abstractions $(\phi(S),A,R)$ might not maintain the MDP structure. This complicates the process of learning $Q_{\phi}^{\pi}$ and $w_{\phi}^{\pi}$. These challenges motivate us to develop model-based irrelevance conditions in the subsequent section.
@@ 3.2 Model-based irrelevance conditions
    The second perspective offers a backward view by reversing the time order. Specifically, due to the symmetric nature of the Markov assumption — implying that if the future is independent of the past given the present, the past must also be independent of the future given the present — the reversed state-action pairs also maintain the Markov property. Leveraging this property, we define another backward MDP, which forms the basis for deriving model-based conditions for achieving $w^{\pi}$-irrelevance and motivates the subsequent two-step procedure. This development represents one of our main contributions.
Figure 2: Illustrations of (a) the forward MDP model and (b) the backward MDP model.
Forward MDP-based model-irrelevance. We first explore the relationship between the model-irrelevance given in Definition 3, and the notions of $Q^{\pi}$-, $\rho^{\pi}$- and $w^{\pi}$-irrelevance.
Theorem 2 (OPE under model-irrelevance)
@@
Let $\phi$ denote a model-irrelevant abstraction.
  • If $\phi$ is additionally $\pi$-irrelevant, then $\phi$ is also $Q^{\pi}$-irrelevant.
  • While $\phi$ is not necessarily $w^{\pi}$-irrelevant, MIS (Method 3) remains valid when applied to the abstract state space. Indeed, the validity only requires reward-irrelevance (see the first part of (2)).
@@
  • While $\phi$ is not necessarily $\rho^{\pi}$-irrelevant, SIS (Method 2) remains valid when applied to the abstract state space if $\phi$ is additionally $\pi$-irrelevant.
@@
  • DRL (Method 4) remains valid when applied to the abstract state space.
@@
The first bullet point establishes the link between model-irrelevance and $Q^{\pi}$-irrelevance, thus proving the validity of the Q-function-based method when applied to the abstract state space. To satisfy $Q^{\pi}$-irrelevance, we need both model-irrelevance and $\pi$-irrelevance. In our implementation, we first adapt existing algorithms (Ha & Schmidhuber, 2018; François-Lavet et al., 2019; Gelada et al., 2019) to train a model-irrelevant abstraction $\phi$, parameterized via deep neural networks. We next combine $\phi(s)$ with $\{\pi(a|s):a\in\mathcal{A}\}$ to obtain a new abstraction $\phi_{for}(s)$. This augmentation ensures $\phi_{for}(s)$ is $\pi$-irrelevant, and hence $Q^{\pi}$-irrelevant. Refer to Appendix B.1 for the detailed procedures.
@@
The last three bullet points prove the validity of the SIS, MIS and DRL, despite $\phi$ being neither $w^{\pi}$-irrelevant nor $\rho^{\pi}$-irrelevant. By definition, $\rho^{\pi}$-irrelevance can be achieved by selecting state features that adequately predict the IS ratio. However, methods for constructing $w^{\pi}$-irrelevant abstractions remain less clear. In the following, we introduce a backward MDP model-based irrelevance condition that ensures $w^{\pi}$-irrelevance. We also note that findings similar to those in the first two bullet points have previously been documented in Li et al. (2006) and Pavse & Hanna (2023), respectively. However, the properties of SIS and DRL estimators under model-irrelevance conditions as summarized in our last two bullet points, remain unexplored in the existing literature.
@@
Backward MDP-based model-irrelevance. To illustrate the rationale behind the proposed model-based abstraction, we introduce the backward MDP model by reversing the time index. Under the (forward) MDP model assumption described in Section 2.1 and that the behavior policy $b$ is not history-dependent, actions and states following $S_{t}$ are independent of those occurred prior to the realization of $S_{t}$. Accordingly, $(S_{t-1},A_{t-1})$ is conditionally independent of $\{(S_{k},A_{k})\}_{k>t}$ given $S_{t}$. Recall that $T$ corresponds to the termination time of trajectories in $\mathcal{D}$. We define a time-reversed process consisting of state-action-reward triplets $\{(S_{t},A_{t},\rho^{\pi}(A_{t},S_{t})):t=T,\dots,1\}$. Its dynamics is described as follows (see also Figure 2(b) for the configuration):
@@
  • Reward generation: For each state-action pair $(S_{t},A_{t})$, we manually set the reward to the IS ratio $\rho^{\pi}(A_{t},S_{t})$, which plays a crucial role in constructing IS estimators.
Given this MDP, analogous to Definition 3, our objective is to identify a state abstraction that is crucial for predicting the reward (e.g., the IS ratio) and the reversed transition function. We provide the formal definition of the proposed backward MDP-based model-irrelevance (short for backward-model-irrelevance) below.
@@
The conditions of backward-model-irrelevance are similar to those specified for model-irrelevance outlined in Definition 3. The first condition (i) essentially requires reward-irrelevance, i.e., $\rho^{\pi}$-irrelevance, in the backward MDP. The second condition in equation (8) is equivalent to the conditional independence assumption between the pair $(A_{t},\phi(S_{t}))$ and $S_{t+1}$ given $\phi(S_{t+1})$. As previously assumed, $S_{t}$ can be decomposed into the union of relevant features $\phi(S_{t})$ and irrelevant features $\psi(S_{t})$, leading to the following factorization:
@@
This indicates a two-step transition in the forward model: initially from $(\phi(S_{t}),A_{t})$ to $\phi(S_{t+1})$, and then from $\phi(S_{t+1})$ to $\psi(S_{t+1})$. Importantly, the generation of $\psi(S_{t+1})$ in the second step is conditionally independent of $A_{t}$ and $\phi(S_{t})$. Consequently, $\phi$ extracts state representations that are influenced either by past actions or past relevant features; see Figure 1(b) for an illustration. Combined with $\rho^{\pi}$-irrelevance, this ensures that all information contained within the historical IS ratios $\{\rho^{\pi}(A_{k},S_{k})\}_{k<t}$ can be effectively summarized using a single $A_{t-1}$ and the abstract state $\phi(S_{t-1})$, thus achieving $w^{\pi}$-irrelevance (see Theorem 3 below).
@@
Assume $\phi$ is backward-model-irrelevant.
  • $\phi$ is both $\rho^{\pi}$-irrelevant and $w^{\pi}$-irrelevant.
  • While $\phi$ is not necessarily $Q^{\pi}$-irrelevant, the Q-function-based method (Method 1) remains valid when applied to the abstract state space.
  • DRL (Method 4) remains valid when applied to the abstract state space.
The first bullet point in Theorem 3 validates the two IS methods when applied to the abstract state space under the proposed backward-model-irrelevance, whereas the last two bullet points validate the Q-function-based method and DRL.
To conclude this section, we draw a connection between the proposed backward-model-irrelevant abstraction for OPE and the Markov state abstraction (MSA) developed by Allen et al. (2021) for policy learning. MSA impose two conditions: (i) inverse-model-irrelevance, which requires $A_{t}$ to be conditionally independent of $S_{t}$ and $S_{t+1}$ given $\phi(S_{t})$ and $\phi(S_{t+1})$; (ii) density-ratio-irrelevance, which requires $\phi(S_{t})$ to be conditionally independent of $S_{t+1}$ given $\phi(S_{t+1})$. For effective policy learning, MSA requires both conditions to hold in data generating processes following a diverse range of behavior policies. When restricting them to one behavior policy, the two conditions are closely related to our backward-model-irrelevance. In particular, they imply our proposed condition in (8) whereas (8) in turn yields density-ratio-irrelevance. This allows us to adapt their algorithm to train state abstractions that satisfy backward-model-irrelevance; see Appendix B.2 for details.
Figure 3: Illustrations of (a) the two-step procedure and (b) an MDP with three groups of state variables, denoted by $\{S_{t,1}\}_{t}$, $\{S_{t,2}\}_{t}$ and $\{S_{t,3}\}_{t}$.
3.3 Two-step procedure for forward and backward state abstraction
The proposed two-step procedure proceeds as follows (see Figure 3(a) for a visualization):
@@ 3.3 Two-step procedure for forward and backward state abstraction
    Backward abstraction: Learn an abstraction $\phi_{2}$ from the abstract state space $\mathcal{X}_{1}$ to $\mathcal{X}_{2}$ using the data triplets $(\phi_{1}(S),A,R)$ that is backward-model-irrelevant.
@@
    Output $\mathcal{X}_{2}$ for off-policy evaluation.
To summarize, our approach sequentially applies the forward and backward abstraction on the state obtained from the previous iteration, progressively reducing state cardinality. To elaborate the usefulness of the two-step procedure in reducing state cardinality, we first analyze a toy example.
@@
A toy example: Consider an MDP where the state variables can be classified into three groups, depicted in Figure 3(b). For this example, we focus on a specific type of state abstraction known as variable selection, which selects a sub-vector from the original state. Key observations from this example are as follows: (i) The reward depends on the state only through the first group of variables; (ii) The evolution of the first group of variables depends only on the second group, and this dependency is indirect. Specifically, the second group evolves first at each time step and subsequently influences the first group; (iii) The second and third groups in the MDP evolve independently, each relying solely on their own previous states; (iv) The behavior policy depends only on the last two groups; (v) Only the second group of variables is directly influenced by the previous action.
According to (i), selecting the first group of variables achieves reward-irrelevance. Combined with (ii) and (iii), choosing the first two groups achieves model-irrelevance. Assuming the target policy is agnostic to the state, the proposed forward abstraction will select the first two groups of variables.
According to (iv) and that the target policy is state-agnostic, selecting the last two groups attains $\rho^{\pi}$-irrelevance. Meanwhile, according to (ii) and (v), selecting these variables also achieves backward-model-irrelevance. Thus, the proposed backward abstraction will select the last two groups.
In the two-step procedure, the forward abstraction first eliminates the third group of variables. Given conditions (ii)-(v), selecting just the second group suffices to achieve backward-model-irrelevance, leading to the elimination of the first group in the subsequent backward abstraction. After two iterations, the procedure produces only one group of variables, demonstrating its efficiency in reducing dimensions compared to using either forward or backward abstraction alone.
In more complex scenarios, each abstraction guarantees that the cardinality of the state space does not increase, effectively maintaining or reducing complexity. The reduction is more likely because forward and backward abstractions, as illustrated in Figures 1(a) and (b), differ by definition. Meanwhile, according to Theorems 2 and 3, the post-abstraction-OPE remains valid for any of the four methods.
Theorem 4 (The two-step procedure)
The four OPE methods remain valid when applied to the abstracted state produced by the proposed two-step procedure.
Finally, we note that one may further consider an iterative procedure that alternates between forward and backward abstractions. However, it remains unclear whether these methods have guarantees.

### v2

@@ 4 Proposed state abstractions for policy evaluation
This section presents model-free (Section 4.1), model-based irrelevance conditions (Section 4.2) for OPE and analyzes the OPE estimators under these conditions (Lemma 1 & Theorem 1). Motivated by this analysis, we propose our iterative procedure (Section 4.3) and study its property (Theorem 2).
Our theoretical analysis is concerned with the Fisher consistency of various OPE estimators, named after, Ronald Fisher, the founder of modern statistics. In particular, the Fisher consistency requires an estimator to be exactly equal to the ground truth given infinite samples. Specialized to our settings, it imposes two requirements:
  1. (i)
    The identification formulas presented in Section 3.3 remain valid when replacing the oracle Q-function or (M)IS ratio with those projected into the proposed abstract state space;
  2. (ii)
    The Q-function or (M)IS ratio defined on the abstract state space is identifiable.
Once the Fisher consistency is established, the estimator’s finite sample properties can be readily obtained using existing techniques (see e.g., Uehara et al., 2021a, ). Therefore, for conciseness and to avoid redundancy, we chose not to present finite sample results in our theoretical analysis.
@@ 4.1 Model-free irrelevance conditions
These irrelevance conditions encourage us to conduct OPE on the abstract state space to reduce sample complexity. Nevertheless, methods for deriving abstractions that satisfy these conditions (particularly $Q^{\pi}$- and $w^{\pi}$-irrelevance) remain unclear. Furthermore, the state-action-reward triplets transformed via these abstractions $(\phi(S),A,R)$ might not maintain the MDP structure. This complicates the process of learning $Q_{\phi}^{\pi}$ and $w_{\phi}^{\pi}$. These challenges motivate us to consider model-based irrelevance conditions introduced in the subsequent section.
@@ 4.2 Model-based irrelevance conditions
    The second perspective offers a backward view by reversing the time order. Specifically, he symmetric nature of the Markov assumption — implying that if the future is independent of the past given the present, the past must also be independent of the future given the present — the reversed state-action pairs also maintain the Markov property. Leveraging this property, we define another backward MDP, which forms the basis for deriving model-based conditions for achieving $w^{\pi}$-irrelevance and motivates the subsequent iterative procedure. This development represents one of our main contributions.
Figure 2: Illustrations of (a) the forward MDP model and (b) the backward MDP model. $b_{t}$ is a shorthand for $b(A_{t}|S_{t})$ for any $t\geq 1$.
Forward MDP-based model-irrelevance. We first discuss the connections between the model-irrelevance given in Definition 3 and the notions of $Q^{\pi}$-, $\rho^{\pi}$- and $w^{\pi}$-irrelevance, and introduce the following conditions to establish Fisher consistency. These conditions are mild and frequently imposed in the RL and OPE literature (see e.g., Thomas and Brunskill,, 2016; Kallus and Zhou,, 2018; Chen and Jiang,, 2019; Fan et al.,, 2020; Cai et al.,, 2021; Shi et al.,, 2021; Kallus and Uehara,, 2022).
Assumption 1 (Boundedness)
All immediate rewards are uniformly bounded.
Assumption 2 (Coverage)
The denominator in (2) is strictly positive.
Assumption 3 (Stationary)
The MDP $(S_{t},A_{t},R_{t})_{t\geq 1}$ is stationary over time.
The following lemma summarizes the findings. Results in the first two bullet points are based on those in the existing literature (see e.g., Li et al.,, 2006; Pavse and Hanna,, 2023).
Lemma 1 (OPE under model-irrelevance)
@@
Assume Assumptions 13 hold. Let $\phi$ denote a model-irrelevant abstraction. Suppose $\phi$ is additionally $\pi$-irrelevant. Then:
  • $Q^{\pi}$-irrelevance & Fisher consistency of Q-function-based method: $\phi$ is also $Q^{\pi}$-irrelevant, and the corresponding Q-function-based method (Method 1) is thus Fisher consistent, i.e., the Q-function $Q_{\phi}^{\pi}$ defined on the abstract space is identifiable and satisfies $\mathbb{E}[f_{1}(Q^{\pi})]=\mathbb{E}[f_{1}(Q^{\pi}_{\phi})]$;
@@
  • Fisher consistency of MIS: While $\phi$ is not necessarily $w^{\pi}$-irrelevant, MIS (Method 3) is Fisher consistent when applied to the abstract state space, i.e., the MIS ratio $w_{\phi}^{\pi}$ defined on the abstract state space is identifiable and satisfies $\mathbb{E}[f_{3}(w^{\pi})]=\mathbb{E}[f_{3}(w^{\pi}_{\phi})]$;
  • Fisher consistency of SIS: While $\phi$ is not necessarily $\rho^{\pi}$-irrelevant, SIS (Method 2) with a history-dependent IS ratio (detailed in the proof of Lemma 1 in Appendix D.1) is Fisher consistent when applied to the abstract space;
@@
  • Fisher consistency of DRL: DRL (Method 4) is Fisher consistent when applied to the abstract state space.
@@
The first bullet point establishes the link between model-irrelevance and $Q^{\pi}$-irrelevance and proves the Fisher consistency of the Q-function-based method when applied to the abstract state space. To satisfy $Q^{\pi}$-irrelevance, we need both model-irrelevance and $\pi$-irrelevance. In our implementation, we first adapt existing algorithms to train a model-irrelevant abstraction $\phi$, parameterized via deep neural networks. We next combine $\phi(s)$ with $\{\pi(a|s):a\in\mathcal{A}\}$ to obtain a new abstraction $\phi_{for}(s)$. This augmentation ensures $\phi_{for}(s)$ is $\pi$-irrelevant, and hence $Q^{\pi}$-irrelevant. Refer to Appendix B.1 for the detailed procedures.
@@
The last three bullet points validate the SIS, MIS and DRL estimators, despite $\phi$ being neither $w^{\pi}$-irrelevant nor $\rho^{\pi}$-irrelevant. By definition, $\rho^{\pi}$-irrelevance can be achieved by selecting state features that adequately predict the IS ratio. However, methods for constructing $w^{\pi}$-irrelevant abstractions remain less clear. In the following, we introduce a backward MDP model-based irrelevance condition that ensures $w^{\pi}$-irrelevance.
@@
Backward MDP-based model-irrelevance. To illustrate the rationale behind the proposed model-based abstraction, we introduce the backward MDP model by reversing the time index. Under the (forward) MDP model assumption described in Section 3.1 and that the behavior policy $b$ is not history-dependent, actions and states following $S_{t}$ are independent of those occurred prior to the realization of $S_{t}$. Accordingly, $(S_{t-1},A_{t-1})$ is conditionally independent of $\{(S_{k},A_{k})\}_{k>t}$ given $S_{t}$. Recall that $T$ corresponds to the termination time of trajectories in $\mathcal{D}$. We define a time-reversed process consisting of state-action-reward triplets $\{(S_{t},A_{t},b(A_{t}|S_{t})):t=T,\dots,1\}$. Its dynamics is described as follows (see also Figure 2(b) for the configuration):
@@
  • Reward generation: For each state-action pair $(S_{t},A_{t})$, we manually set the reward to the behavior policy $b(A_{t}|S_{t})$, which plays a crucial role in constructing IS estimators.
Given this MDP, analogous to Definition 3, our objective is to identify a state abstraction that is crucial for predicting the reward (e.g., the behavior policy) and the reversed transition function. We provide the formal definition of the proposed backward MDP-based model-irrelevance (short for backward-model-irrelevance) below.
@@
The conditions of backward-model-irrelevance are similar to those specified for model-irrelevance outlined in Definition 3. Equation (4) essentially requires behavior-policy-irrelevance, or reward-irrelevance in the backward MDP. Equation (5) corresponds to the “backward-transition-irrelevance”, and is equivalent to the conditional independence assumption between the pair $(A_{t},\phi(S_{t}))$ and $S_{t+1}$ given $\phi(S_{t+1})$. As previously assumed, $S_{t}$ can be decomposed into the union of relevant features $\phi(S_{t})$ and irrelevant features $\psi(S_{t})$ (see Figure 1), leading to the following factorization:
@@
This indicates a two-step transition in the forward model: initially from $(\phi(S_{t}),A_{t})$ to $\phi(S_{t+1})$, and then from $\phi(S_{t+1})$ to $\psi(S_{t+1})$. Importantly, the generation of $\psi(S_{t+1})$ in the second step is conditionally independent of $A_{t}$ and $\phi(S_{t})$. Consequently, $\phi$ extracts state representations that are influenced either by past actions or past relevant features; see Figure 1(b) for an illustration. Combined with $\pi$-irrelevance and behavior-policy-irrelevance, this ensures that all information contained within the historical IS ratios $\{\rho^{\pi}(A_{k},S_{k})\}_{k<t}$ can be effectively summarized using a single $A_{t-1}$ and the abstract state $\phi(S_{t-1})$, thus achieving $w^{\pi}$-irrelevance (see Theorem 1 below).
@@
Assume Assumptions 13 hold, and $\phi$ is both backward-model-irrelevant and $\pi$-irrelevant.
  • Then $\phi$ is both $\rho^{\pi}$-irrelevant and $w^{\pi}$-irrelevant.
  • Additionally, all the four OPE methods (i.e., Q-function-based, SIS, MIS, DRL) are Fisher consistent when applied to the abstract state space.
Theorem 1 validates the four OPE methods when applied to the abstract state space. To conclude this section, we draw a connection between the proposed backward-model-irrelevant abstraction for OPE and the Markov state abstraction (MSA) developed by Allen et al., (2021) for policy learning. MSA impose two conditions: (i) inverse-model-irrelevance, which requires $A_{t}$ to be conditionally independent of $S_{t}$ and $S_{t+1}$ given $\phi(S_{t})$ and $\phi(S_{t+1})$; (ii) density-ratio-irrelevance, which requires $\phi(S_{t})$ to be conditionally independent of $S_{t+1}$ given $\phi(S_{t+1})$. For effective policy learning, MSA requires both conditions to hold in data generating processes following a diverse range of behavior policies. When restricting them to one behavior policy, the two conditions are closely related to our backward-model-irrelevance. In particular, they imply our proposed backward-transition-irrelevance condition in (5) whereas backward-transition-irrelevance in turn yields their density-ratio-irrelevance. This allows us to adapt their algorithm to train state abstractions that satisfy our proposed backward-model-irrelevance; see Appendix B.2 for details.
Finally, it is worth noting that both the proposed backward-model-irrelevant abstraction and MSA require the behavior policy to be Markov, independent of the past observations.The following section will relax this condition and extend our proposal to accommodate history-dependent behavior policies.
Figure 3: Illustrations of the iterative procedure.
4.3 Iterative procedure for deep state abstraction
To summarize, we have reviewed the model-irrelevance condition and proposed a new backward-model-irrelevance condition. Both lead to Fisher consistent OPE estimators when confined to the corresponding abstract state spaces. This motivates us to combine the two procedures for a more condensed state abstraction, resulting in the following iterative algorithm (see Figure 3 for a visualization):
@@ 4.3 Iterative procedure for deep state abstraction
    Backward abstraction: Learn an abstraction $\phi_{2}$ from the abstract state space $\mathcal{X}_{1}$ to $\mathcal{X}_{2}$ using the data triplets $(\phi_{1}(S),A,R)$ that is both backward-model-irrelevant and $\pi$-irrelevant.
@@
    Iterate the two steps to compute the final abstraction $\phi_{K}\circ\cdots\circ\phi_{2}\circ\phi_{1}$ from the ground space $\mathcal{S}$ to $\mathcal{X}_{K}$ where $K$ denotes the number of iterations and $\circ$ denotes the function composition operator.
In particular, our approach alternates between forward and backward abstraction on the state space obtained from the previous iteration. Each iteration guarantees that the cardinality of the state space does not increase, effectively maintaining or reducing complexity. Consequently, such an iterative procedure progressively reducing state cardinality, which ultimately yields a deeply-abstracted state. We thus refer to our approach as deep state abstraction (DSA).
To elaborate the usefulness of DSA in reducing state cardinality, we analyze two examples: a bandit example and an MDP example. In both examples, we focus on a specific type of state abstraction known as variable selection, which selects a sub-vector from the original state. Additionally, we focus on the class of state-agnostic target policies where $\pi$ is independent of the states. This type of policy is prevalent in causal inference and A/B testing, where the objective is to learn the global treatment effect of applying either a new or old policy consistently over time (see e.g., Hu and Wager,, 2022; Leung,, 2022; Bojinov et al.,, 2023; Shi et al.,, 2023; Xiong et al.,, 2024).
Figure 4: Two illustrative examples.
Example I (A bandit example): To illustrate the main idea, we start by considering the contextual bandit setting (CB, Dudík et al.,, 2014), which can be regarded as a special MDP with independent state transitions. Under this setting, the states are i.i.d. generated, and model-irrelevance is reduced to reward-irrelevance whereas the proposed backward-model-irrelevance simplifies to behavior-policy-irrelevance. When specialized to variable selection in CB, our proposal is reduced to the iterative confounder selection algorithm in causal inference (see e.g., Guo et al.,, 2022); see also the review of confounder selection in Appendix A. We assume the states can be divided into three independent groups, denoted by $S_{t}^{(1)}$, $S_{t}^{(2)}$ and $S_{t}^{(3)}$, respectively. Each group influences the system differently: $S_{t}^{(1)}$ affects only the reward, $S_{t}^{(2)}$ impacts both the action and the reward, and $S_{t}^{(3)}$ solely influences the action; see Figure 4(a) for an illustration. As formally proven in Lemma D.3 (see Appendix D.5):
  • The forward abstraction selects the first two groups $S_{t}^{(1)}$ and $S_{t}^{(2)}$;
  • The proposed backward abstraction selects the last two groups $S_{t}^{(2)}$ and $S_{t}^{(3)}$;
  • The proposed DSA selects their intersection $S_{t}^{(2)}$ and converges in two steps, resulting in a smaller subset of variables compared to the two non-iterative procedures.
@@
Example II (An MDP example): We next consider an MDP with three groups of states, depicted in Figure 4(b). Key observations from this example are as follows: (i) The reward depends on the state only through the first group of variables; (ii) The evolution of the state depends only on the second group. Specifically, the second group evolves first at each time step and subsequently influences the rest two groups; (iii) The behavior policy depends only on the last group; (iv) The second group is directly influenced by the previous action. Based on these observations, we show have that:
  • According to (i), selecting the first group of variables achieves reward-irrelevance.
  • According to (iii), selecting the last group of variables achieves behavior-policy-irrelevance.
  • According to (ii) and (iv), selecting the second group achieves both transition-irrelevance (see the second equation in 3) and backward-transition-irrelevance (see (5)).
Consequently, the forward and backward abstractions select the first two and last two groups of variables, respectively. The iterative procedure again selects their intersection $S_{t}^{(2)}$ and converges in two iterations, leading to in a smaller state space. The reader is referred to Appendix D.5 for formal justifications.
In both examples, we have demonstrated the advantage of DSA in reducing state space cardinality over non-iterative procedures. However, in more general scenarios, two challenges arise from the iterative nature of DSA: (i) After forward abstraction, the behavior policy when restricted to the abstract space can be history-dependent. This would invalidate the subsequent backward abstraction for achieving $w^{\pi}$-irrelevance. (ii) After backward abstraction, the Markov assumption might be violated. This would invalidate the subsequent forward abstraction for consistent OPE.
To address both challenges, we modify the proposed backward-model-irrelevance by employing a history-dependent behavior policy $b_{t}(a_{t}|s_{t},a_{t-1},s_{t-1},\cdots,s_{1})=\mathbb{P}(A_{t}=a_{t}|S_{t}=s_{t},A_{t-1}=a_{t-1},S_{t-1}=s_{t-1},\cdots,S_{1}=s_{1})$ in (4). Specifically, we require that for any two state sequences $\{s_{\ell}^{(1)}\}_{\ell}$ ,$\{s_{\ell}^{(2)}\}_{\ell}$ such that $\phi(s_{\ell}^{(1)})=\phi(s_{\ell}^{(2)})$, any action sequence $\{a_{\ell}\}_{\ell}$ and any time index $t\geq 1$,
$b_{t}(a_{t}|s_{t}^{(1)},a_{t-1},s_{t-1}^{(1)},\cdots,s_{1}^{(1)})=b_{t}(a_{t}|s_{t}^{(2)},a_{t-1},s_{t-1}^{(2)},\cdots,s_{1}^{(2)}).$ (6)
The following theorem validates the abstraction produced by the proposed DSA at any iteration.
Theorem 2 (OPE under the iterative procedure)
Assume Assumptions 13 hold. With the refined backward-model-irrelevance, the four OPE methods are Fisher consistent when applied to the abstract state space produced by the proposed DSA, regardless of the number of iterations conducted.
Finally, we note that the initialization of the iterative procedure doesn’t necessarily have to begin with forward abstraction; backward abstraction can serve as the starting point as well. In our experiments, both starting points have their merits, with their effectiveness varying depending on the environment. However, the overall differences in results are small.